OpenStudy (anonymous):

If the sides of a square are increased by 3cm, its area becomes 144cm^2. Find the length of the sides of the original square.

OpenStudy (anonymous):

let the side be x (x + 3)^2 = 144

OpenStudy (anonymous):

O.O WAIToops wronggg question SORRY change to this: If the sides of a square are increased by 4cm, its area becomes 196cm^2. Find the length of the sides of the original square.

OpenStudy (anonymous):

and explain the steps too thanks :D

OpenStudy (ash2326):

Let the lengths of the sides of the original square be x cm it's given that they are increased by 4 and then the area becomes 196 cm^2 so the sides become x+4 Area of square= \(side^2\) so \[(x+4)^2=196\] Do you get this?

OpenStudy (ash2326):

@Steeben ?

OpenStudy (anonymous):

haha yups :D

OpenStudy (ash2326):

could you solve for x?

OpenStudy (anonymous):

oh! 10

OpenStudy (ash2326):

yes, you're right.:)